OpenStudy (anonymous):

AP Calculus: Slope Fields/ Separate Variables Find the equation of the curve, given the derivative and the indicated point on the curve. dy/dx=x

OpenStudy (anonymous):

given point on graph is (0,1/2) @SithsAndGiggles

OpenStudy (anonymous):

\[\frac{dy}{dx}=x\\ dy=x\;dx\\ \int dy=\int x\;dx\] Just like the other problems.

OpenStudy (anonymous):

ok so i use power rule as well to integrate x?

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

is this the equation? \[y=\frac{ 1 }{ x } + C\]

OpenStudy (anonymous):

No. It looks like you should practice some basic integration techniques, specifically the power rule. For the integral \(\int x\;dx,\) you get \(\displaystyle\frac{x^{1+1}}{1+1}+C=\frac{1}{2}x^2+C.\)

OpenStudy (anonymous):

so we use the power rule for this problem as well?

OpenStudy (anonymous):

OHHH wow i'm thinking too much :(

OpenStudy (anonymous):

so is the answer \[y=\frac{ x^2 }{ 2 } + \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Yep, that's correct.

OpenStudy (anonymous):

ok ty :)

OpenStudy (anonymous):

i mite have more questions so.. yeah :)

OpenStudy (anonymous):

but i think i pretty much get it xD

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

@SithsAndGiggles can you help me with another please? :)

OpenStudy (anonymous):

sure can